The length of a rectangle is 8 more than 5 times its width. The perimeter of the rectangle is 544. Find the length of the rectangle.






Answer by Flannery(124)

2(5w+8)+2w=544
10w+16+2w=544
12w+16=544
12w=528
w=44
2(5(44)+8)+2(44)=544
2((220)+8)+88=544
2(228)+88=544
456+88=544
l=p-88/2
l=(544-88)/2
l=456/2
l=228
2(228)+2(44)=544
456+88=544
Length is 228, first we solved for width which is 44.